OpenStudy (anonymous):

A cone has a radius of 12 inches and a surface area of 1050 square inches. find the slant height of the cone to the nearest tenth of an inch. can you also explain the steps?

OpenStudy (bahrom7893):



OpenStudy (anonymous):

Surface area -> convert to height using radius given. Then, use right triangle properties to find slant height (hypotenuse).

OpenStudy (bahrom7893):

http://0.tqn.com/d/math/1/0/t/L/conerr.jpg 1050 = pi*r*x+pi*r^2 1050 = pi*12*x + pi*144

OpenStudy (bahrom7893):

12pix = 1050-144pi

OpenStudy (bahrom7893):

x = (1050-144pi)/(12pi)

OpenStudy (bahrom7893):

just calculate that and round

OpenStudy (anonymous):

thanks so much more help than my online teacher lol